OpenStudy (anonymous):

how do i solve this??? d - 10 = -30

OpenStudy (anonymous):

d - 10 = - 30 d = - 30 + 10 d = -20

OpenStudy (thesecret20111):

Well do you know what a pronumeral is?

OpenStudy (anonymous):

no

OpenStudy (thesecret20111):

It's basically the backbone to all algebra. Its a letter that actually represents a number. so d in this equation is actually representing an unknow number.

OpenStudy (thesecret20111):

So what you're equation d - 10 = -30 is saying is that some unknow number -10 = -30

OpenStudy (thesecret20111):

You have to find out what number would = -30 when you take 10 away from it.

OpenStudy (thesecret20111):

You don't just have to work this out in your head though.. That's why algebra is helpful. You can work this out by using the rule that whatever you do to one side of the equal sign you do to the other as well.

OpenStudy (thesecret20111):

You do this to try to make the letter by itself. So in the equation d - 10 = -30 You want to get rid of the -10 on the side of the '=' sign with the d

OpenStudy (thesecret20111):

To do this you want to make -10 = 0. To do that you would have add 10 it. Don't just do it to that side otherwise you will get the wrong answer.. You have to do the same to both sides of the "=" sign

OpenStudy (thesecret20111):

So you get d - 10+10 = -30+10

OpenStudy (thesecret20111):

Now work out the new numbers: -10+10=0 and -30+10=-20 So that gives us d-0=-20

OpenStudy (thesecret20111):

d-0=-20 is the same thing as d=-20 Meaning that you have found that d represented -20 and that's how simple algebra can be used.